The volume of a sample of hydrogen is 798 mL and it exerts 621 mm Hg pressure at 5.00℃. What volume does it occupy at standard temperature and pressure?
1
Expert's answer
2013-03-11T10:10:12-0400
P1V1/T1 = P2V2/T2
If one of these values is not given then it is considered a constant you just leave that variable out of the equation.
PV=nRT is used when you are working the # of moles or grams. n is the # of moles so if given grams you must convert to moles before you begin.
STP (standard temperature and pressure) P= 1atm ,T= 273K, R= 0.08206 L-atm / mol-K
P1V1/T1 = P2V2/T2
760*V1/273 = 621*798/278
V1*760*278 = 273*621*798
V1 = 640mL